APPENDIX F: Sample TEX Landscape-Mode Overhead Made Using PostScript Fonts Because of the frequent interest in generating overheads for teaching and dissertation-defense purposes, this appendix consists of output for a sample overhead built using TEX. In the context of showing how to make overheads, this example also illustrates how to obtain landscape mode output and how to use native PostScript fonts (such as Helvetica and Helvetica Bold). A machine readable copy of the raw TEX commands used to typeset that resume is available to OREGON VAX users as:

      TEX$DISK:[TEX.SAMPLES]OVERHEAD2.TEX

On other systems get a copy of that file via anonymous FTP.

FTP to DECOY.UOREGON.EDU (128.223.32.19) and see directory pub/tex/samples.